Circular No. 8845

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2006 A8 AND C/2006 Y10–Y17 (SOHO) Nine additional near-sun comets have been found on 2006 SOHO web-

site images (cf. IAUC 8844), all being Kreutz sungrazers except apparently for C/2006 Y12. C/2006 A8 was stellar in appearance and very faint (mag ≈ 7–7.5) with no tail. All of the remaining objects that were visible in C3 images appeared therein as small and stellar; all of the remaining objects that were visible in C2 images appeared therein as tailless and slightly diffuse, except for C/2006 Y12, which was faint yet more stellar in appearance and unusually small. In the table below, new finder ‘AW’ = A. Watson.

Comet  2006 UT  α_{2000}  δ_{2000} Inst. F  MPEC
C/2006 A8 Jan. 10.013 19^{h}36.^{m}4 −24^{o}04′ C3  BZ 2007-L16
C/2006 Y10 Dec. 23.071 18 11.9 −26 33 C3/2 TH 2007-K45
C/2006 Y11  24.821 18 18.4 −25 34 C3/2 MA 2007-K45
C/2006 Y12 27.814 18 31.4 −24 40 C2 RK 2007-K45
C/2006 Y13  28.154 18 35.4 −25 53 C3/2 WX 2007-K45
C/2006 Y14 28.504 18 34.6 −24 38 C2 RK 2007-K45
C/2006 Y15  29.179 18 39.4 −25 51 C3/2 WX 2007-K46
C/2006 Y16  29.413 18 38.6 −24 44 C2  AW 2007-K46
C/2006 Y17  29.529 18 43.9 −25 44 C3  HS 2007-K46

V1280 SCORPII R. C. Puetter, Center for Astrophysics and Space Science, University

of California at San Diego; R. J. Rudy, D. K. Lynch, R. W. Russell, S. Mazuk, and R. L. Pearson, The Aerospace Corporation; C. E. Woodward, University of Minnesota; and R. B. Perry, Langley Research Center, NASA, report 0.4- to 2.5-μm spectroscopy of this nova (cf. IAUC 8803, 8807, 8809, 8812) using the Lick 3-m telescope (+ VNIRIS) on May 6 UT, as well as 0.8to 5.5-μm spectroscopy using the Infrared Telescope Facility 3-m telescope (+ SpeX) on May 31. This nova is still in a very low excitation state, showing strong C I lines and no discernable He I emission. The spectrum shows a significantly narrower [N I] line at 1.04 μm than the permitted emission lines. The reddening derived from the O I lines, which is due in part to the dust shell, is E(B − V) = 1.7.

2007 June 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green